How many kilograms of pure alcohol must be added to 40 kg of a 70% alcohol solution to produce an 85% alcohol solution?
Answer by alicealc(293)

Let pure alcohol = x kg

x + (70/100)*40 = (85/100)*(x + 40)
<=> x + 0.7*40 = 0.85*(x + 40)
<=> x + 28 = 0.85x + 34
<=> x - 0.85x = 34 - 28
<=> 0.15x = 6
<=> x = 6/0.15
<=> x = 40

so 40 kg of pure alcohol must be added